### v1

# Universal Entanglement Revival of Topological Origin

Abstract
We have analyzed the dynamics of entanglement in dissipative fermionic and bosonic Su–Schrieffer–Heeger (SSH) models and found that, when the decoherence channel preserves the chiral symmetry, they exhibit a revival of entanglement in a wide range of parameters. This behavior only emerges in the topological phase, with the visibility of the revival dropping to zero at the phase boundary. Furthermore, the revival acquires a universal character once the system size exceeds the localization length of the edge modes. Our findings indicate that the universal entanglement revival has its origin in the topological properties of the SSH model. These dynamical properties may be experimentally accessible, for example, using photonic quantum computers.
1 Introduction
Entanglement is a critical resource in quantum information processing, yet is intrinsically fragile with respect to interactions with the environment. These induce decoherence, which typically leads to the rapid degradation of entanglement. Nevertheless, certain systems have been shown to exhibit the interesting phenomenon of entanglement revival, in which, after the initial decay, an unexpected recovery occurs at a later stage of the dynamical evolution [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. A robust mechanism causing this type of dynamics has been recently discussed for an ensemble of three-level systems interacting with single-mode bosons [16], where the revival is associated to the presence of a special zero-energy state weakly affected by decoherence. The dissipative evolution rapidly destroys the initial entanglement but eventually turns itself into a self-purification process, resulting in a distinct entanglement revival.
The mechanism of Ref. [16] is generic, thus might be applicable to other scenarios. Indeed, we show here that topological system are a natural candidate to observe this type of revivals. The potential relevance of Ref. [16] for topological systems is suggested by the crucial role played in the three-level ensemble by a chiral symmetry, and by the fact that the state responsible for the revival belongs to the zero-energy subspace associated with that symmetry. In this sense, the metastable state of Ref. [16] is analogous to the zero-energy edge mode of a topological system.
To explore concretely these ideas, we specialize our analysis to a dissipative version of the Su-Schrieffer-Heeger (SSH) model. The SSH model is a well-known paradigm for topological insulators [17, 18, 19] and has considerable relevance for quantum information and quantum computing applications [20, 21, 22, 23, 24, 25], due to its robustness against disorder and defects [26, 27, 28, 29, 30, 31, 32]. Recent studies have increasingly focused on how topological systems, including the SSH model, behave under dissipative dynamics [33, 34, 35, 36, 37, 38]. Gaining a deeper understanding of dissipation in topologically protected systems is critical for furthering the practical implementation of quantum technologies. Here we explore how the topological properties of the SSH model influence entanglement dynamics, which is of significant interest for the development of more resilient quantum systems.
The topological edge states of the SSH model remain stable when decoherence only acts on one of the two sublattices (specifically, the odd sites). Under this scenario, we find the occurrence of an entanglement revival directly tied to the presence of a topological edge mode. We also show that the revival approaches a universal form when increasing the system size, i.e., the evolution of entanglement becomes independent of the chain length. Therefore, the presence (absence) of the revival is a definite signature of the topological (trivial) phase. Our findings apply to both fermionic and bosonic versions of the SSH model, thus they could be simulated in a controlled manner using photonic quantum computers [39, 40, 41]. Nearest-neighbor hoppings of variable coupling strengths can be realized with tunable beam splitters. Additionally, the required photon loss pattern — lower at even sites and higher at odd sites — can be simulated by selective leakage of photons. Therefore, we expect that the topologically protected entanglement revivals discussed here are achievable with current photonic technology.
This paper is organized as follows: In Section 2, we introduce the SSH model, briefly reviewing its chiral symmetry and topological properties. The theoretical methods we used are detailed in Section 3, including the semiclassical approach and different entanglement measures. Section 4 presents the results, highlighting the universal entanglement revival observed in both fermionic and bosonic SSH models as well as its connection to topological edge states. Section 5 concludes the paper. To further support the arguments in the main text, two appendices are provided: A demonstrates the universal revival behavior in the mutual information as an alternative extent of entanglement. B examines the density of von Neumann and Rényi entropies, which also exhibit a universal behavior.
2 Model
Figure 1: Schematics of the dissipative Su-Schrieffer-Heeger (SSH) model. The coupling strengths between nearest-neighbor sites alternate between two values, $g_{1}$ and $g_{2}$, and the quantum decoherence process (wavy lines) occurs only at the odd sites. For numerical efficiency, we ignore the shaded region and denote the remaining part as ‘half-SSH chain’.
The SSH model is a prototype of one-dimensional (1D) topological insulators [18, 19]. As a tight-binding model with alternating hopping amplitudes, $g_{1}$ and $g_{2}$, it exhibits a chiral symmetry (also known as sub-lattice symmetry), which, together with the time-reversal symmetry, completely classifies the topological states of matter. The chiral symmetry is characterized by a unitary operator that anti-commutes with the Hamiltonian, and ensures that the Hamiltonian matrix can be written in a block-off-diagonal form, leading to a degenerate zero-energy subspace and a symmetric energy spectrum around zero energy. The zero-energy states are especially interesting: According to the bulk-boundary correspondence principle, they always accompany the topologically non-trivial phase of topological insulators and are localized at the boundaries of the system. In the SSH model, the topological phase occurs when $g_{1}<g_{2}$  [^1] Without loss of generality, we assume real and positive hopping amplitudes. Complex hopping amplitudes can always be brought to real positive values with a gauge transformation.. In this phase, there are two zero-energy edge states at the two ends of the chain. The localization length of each edge state describes how far the state spreads from the edge into the bulk of the chain, and is given by $\xi=2/\log[g_{2}/g_{1}]+1$.
For systems of sufficiently large size $L\gg\xi$, the zero-energy edge states at the two ends are decoupled from each other. As the effects of interest in this work only involve a single edge state, we will, for the sake of numerical efficiency, consider only half of the physically realistic chain, ignoring the shaded portion in Fig. 1  [^2] In this half-chain model one unit cell is not complete, as it contains only a single site instead of two. Therefore, the bulk-boundary correspondence does not hold in the usual sense.. This simplification allows us to concentrate on the properties and behavior of a single edge state, while maintaining the essential features of the topological phase. This half-SSH model contains $n$ two-site unit cells and a single unpaired site, so the overall chain length $L=2n+1$ is always odd. The Hamiltonian reads:
$\hat{H}=\sum_{i=1}^{(L-1)/2}\left(g_{1}\hat{f}_{2i-2}^{\dagger}\hat{f}_{2i-1}+g_{2}\hat{f}_{2i-1}^{\dagger}\hat{f}_{2i}\right)+\mathrm{H.c.},$ (1)
where $\hat{f}_{i}^{\dagger}$ and $\hat{f}_{i}$ are, respectively, the creation and annihilation operators at site $i=0,1,\ldots,L-1$. In the fermionic (bosonic) SSH model, they satisfy the anti-commutation (commutation) relations, $\{\hat{f}_{i},\hat{f}_{j}^{\dagger}\}=\delta_{ij}$ $([\hat{f}_{i},\hat{f}_{j}^{\dagger}]=\delta_{ij}).$ To better reveal the consequences of the chiral symmetry and the topological properties of the system, we rewrite the Hamiltonian in terms of the normal modes $\hat{C}_{\pm k}$:
$\hat{H}=\sum_{k=1}^{(L-1)/2}\epsilon_{k}\left(\hat{C}_{k}^{\dagger}\hat{C}_{k}-\hat{C}_{-k}^{\dagger}\hat{C}_{-k}\right),$ (2)
where the single-particle energies ±$\pm\epsilon_{k}$ are given by:
$\epsilon_{k}:=\sqrt{g_{1}^{2}+g_{2}^{2}+2g_{1}g_{2}\cos(\tilde{k})},$ (3)
with $\tilde{k}:=2\pi k/(L+1).$ The normal modes are related to the site operators $\hat{f}_{i}$ by
$\hat{C}_{\pm k}=\sum_{i=0}^{(L-1)/2}A_{ki}\hat{f}_{2i}\pm\sum_{i=1}^{(L-1)/2}B_{ki}\hat{f}_{2i-1},$ (4)
with
$A_{ki}$ $:=\sqrt{\frac{2}{L+1}}\sin\left(\tilde{k}i+\phi_{k}\right),$ (5a)
$B_{ki}$ $:=\sqrt{\frac{2}{L+1}}\sin\left(\tilde{k}i\right),$ (5b)
and $\cos(\phi_{k}):=(g_{1}\cos\tilde{k}+g_{2})/\epsilon_{k}.$ Furthermore, there is a special normal mode with zero energy, which does not appear in Eq. (2):
$\hat{C}_{0}=\sqrt{\frac{1-(g_{1}/g_{2})^{2}}{1-(g_{1}/g_{2})^{L+1}}}\sum_{i=0}^{(L-1)/2}(-g_{1}/g_{2})^{i}\hat{f}_{2i}.$ (6)
Clearly, when $g_{1}<g_{2}$ this mode has its probability amplitudes localized at the left end of the system  [^3] In the full SSH model, there is another zero-energy edge mode at the right end.. For later use, when $k=0$ we define:
$A_{0i}=\sqrt{\frac{1-(g_{1}/g_{2})^{2}}{1-(g_{1}/g_{2})^{L+1}}}(-g_{1}/g_{2})^{i}.$ (7)
Recall that, unlike this zero-energy edge mode, all other modes $\hat{C}_{\pm k}$ ($k\neq 0$) are bulk states with probability amplitudes spread over the whole chain.
We are interested in the topological influence on the dynamics of quantum entanglement in the presence of decoherence. To clearly illustrate the topological effects, we assume that the quantum decoherence processes preserve the chiral symmetry; that is, we assume that decoherence processes occur only on the odd sites $\hat{f}_{2i-1}$, with a uniform decay rate $\Gamma$. The noisy dynamics is governed by the quantum master equation
$\frac{d\hat{\rho}}{dt}=-i[\hat{H},\hat{\rho}]+\Gamma\sum_{i=1}^{(L-1)/2}\mathcal{D}[\hat{f}_{2i-1}]\hat{\rho},$ (8)
where $\hat{\rho}$ is the density operator of the system and $\mathcal{D}$ is the superoperator $\mathcal{D}[\hat{f}]\hat{\rho}=\hat{f}\hat{\rho}\hat{f}^{\dagger}-(\hat{f}^{\dagger}\hat{f}\hat{\rho}+\hat{\rho}\hat{f}^{\dagger}\hat{f})/2,$ describing the quantum decoherence processes.
We will consider an initial state of the form:
$|\psi_{g}\rangle=\hat{C}_{0}^{\dagger}\Pi_{k=1}^{(L-1)/2}\hat{C}_{-k}^{\dagger}|\rangle,$ (9)
where |$\rangle$| ⟩ is the vacuum state. Since all negative-energy modes and the zero-energy mode are singly occupied, this initial state is one of the two degenerate many-body ground states of the fermionic model. In that case, it is a natural choice for the initial state. In the bosonic model, occupying all negative and zero-energy normal modes can be achieved by using external driving, as commonly done in experimental demonstrations of topological band structures, e.g., in photonic crystals [42, 43]. Since we study non-equilibrium dynamics, the resulting physical effects inevitably depend on the initial state. However, as long as the zero-energy edge mode is initially occupied, our main findings remain qualitatively valid for many other initial states. In particular, we have checked that a universal revival is found for an initial state where all the normal modes are singly occupied. This confirms that the universal entanglement dynamics reported below is closely related to topological features.
3 methods
Before presenting in Sec. 4 our main results, we describe in this Section the methods used to solve the quantum master equation, Eq. (8), and to analyze the entanglement content of the solution.
3.1 Semiclassical approach
The computational cost to exactly solve the quantum master equation increases exponentially with system size  [^4] For non-interacting fermions with a restricted class of quantum jump operators, a polynomial-time method is available. See, for example, Refs. [44, 45].. Fortunately, in the parameter regime of our interest, a semicassical approach provides an accurate description of the dynamics. We first note that, when the coupling strengths are much larger than the decay rate ($g_{1},g_{2}\gg\Gamma$), the eigenstates of Eq. (1) are generally well-separated in energy. Since the coherent phase factor between two eigenstates oscillates rapidly in time, it gets averaged out throughout the slow dissipative evolution and we may assume that $\hat{\rho}(t)$ is approximately diagonal in the many-body eigenstates, $\mathinner{\lvert{\textstyle E_{\alpha}}\rangle}$:
$\hat{\rho}(t)\approx\sum_{\alpha}|E_{\alpha}\rangle P_{\alpha}(t)\langle E_{\alpha}|.$ (10)
The probabilities $P_{\alpha}$ of different many-body eigenstates $|E_{\alpha}\rangle$ satisfy the master equation
$\frac{dP_{\alpha}(t)}{dt}=\sum_{\beta}\left(\gamma_{\alpha\beta}P_{\beta}(t)-\gamma_{\beta\alpha}P_{\alpha}(t)\right),$ (11)
where the transition rate $\gamma_{\alpha\beta}$ from $\mathinner{\lvert{\textstyle E_{\beta}}\rangle}$ to $\mathinner{\lvert{\textstyle E_{\alpha}}\rangle}$ is given by $\gamma_{\alpha\beta}:=\Gamma\sum_{i=1}^{(L-1)/2}|\langle E_{\alpha}|\hat{f}_{2i-1}|E_{\beta}\rangle|^{2}.$ In terms of the eigenmode operators:
$\gamma_{\alpha\beta}:=\frac{\Gamma}{2}\sum_{k\neq 0}|\langle E_{\alpha}|\hat{C}_{k}|E_{\beta}\rangle|^{2},$ (12)
which can be derived using $\sum_{i=1}^{(L-1)/2}B_{ki}B_{k^{\prime}i}=\delta_{kk^{\prime}}/2.$ This relation has its origin in the the chiral symmetry, leading to coefficients $A_{ki}$ and $B_{ki}$, see Eqs. (5) and (7), which are separately normalized. Therefore, we also have $\sum_{i=0}^{(L-1)/2}A_{ki}A_{k^{\prime}i}=\delta_{kk^{\prime}}/2.$
We note some unique properties of the semiclassical dynamics described by Eqs. (10) and (11). First, every time a decay process occurs, the total number of particles decreases by one. However, the zero-energy edge state can never decay, because $\hat{C}_{0}$ has non-zero amplitudes only at the even sites, see Eq. (6). In the long-time limit ($t\to\infty$) only one particle remains, occupying the zero-energy edge state:
$\hat{\rho}_{\infty}:=\lim_{t\rightarrow\infty}\hat{\rho}(t)=\hat{C}_{0}^{\dagger}|\rangle\langle|\hat{C}_{0}.$ (13)
Second, with the initial condition Eq. (9) the occupation numbers of the eigenmodes are either one or zero, and a decay process corresponds to one of the occupied modes becoming empty. From Eq. (12), we see that all these decays occur with an identical rate, $\Gamma/2$, which allows us to solve the semiclassical evolution of Eqs. (10) and (11) in closed form:
$\hat{\rho}(t)\approx\sum_{m=0}^{(L-1)/2}P_{m+1}(t)\sum_{\{k_{i}\}_{i=1}^{m}}\hat{C}_{\{k_{i}\}}^{\dagger}\hat{\rho}_{\infty}\hat{C}_{\{k_{i}\}},$ (14)
with
$P_{m+1}(t)=\exp\left[\frac{-m\Gamma t}{2}\right]\left(1-\exp\left[-\frac{\Gamma t}{2}\right]\right)^{\frac{L-1}{2}-m}.$ (15)
In Eq. (14), we used the notation $\hat{C}_{\{k_{i}\}_{i=1}^{m}}:=\prod_{i=1}^{m}\hat{C}_{-k_{i}}$ where $\{k_{i}\}_{i=1}^{m}$ is an unordered set of indexes (with $k_{i}>0$). We see from Eq. (14) that the probabilities $P_{\alpha}$ of the eigenstates hosting $m-1$ particles with negative energy (and one particle in the $\hat{C}_{0}$ mode) are all equal. For those states we can simply set $P_{\alpha}=P_{m}$.
So far, we have justified the semiclassical approximation in the large-coupling limit $g_{1},g_{2}\gg\Gamma$. However, the above description is more general, and allows us to compute the dynamics deep in the topological phase ($g_{1}\ll g_{2}$), when $g_{1}$ can become comparable to $\Gamma$. In this limit, the energy differences between the eigenstates with the same $m$ appearing in Eq. (14) are of order $g_{1}$, thus the corresponding coherences would not oscillate on a fast timescale. Fortunately, however, the dissipative evolution does not generate such coherences, as can be checked by applying the full master equation, Eq. (8), to a generic eigenstate $|E_{\alpha}\rangle\langle E_{\alpha}|$. This is due to a destructive cancellation of amplitudes, which follows from the orthogonality of the eigenmodes. In particular, $\sum_{i=1}^{(L-1)/2}B_{ki}B_{k^{\prime}i}=0$ when $k\neq k^{\prime}$. The semiclassical approximation is still not exact, as the full dynamics generates some coherence between eigenstates with large energy differences, of order $g_{2}$. Thus, it is still necessary to satisfy the condition $g_{2}\gg\Gamma$.
To demonstrate the accuracy of the semiclassical approximation when $g_{1}\approx\Gamma$, we compare in Fig. 2 two results for the probability $P_{m}(t)$: One from the numerical solution of the full quantum master equation (8) and another from the semiclassical approximation (15). Clearly we observe an excellent agreement, justifying the semiclassical approach. In passing, we note that the maximum probability appears earlier for states with more particles, which is intuitively appealing.
Figure 2: Probability $P_{m}$ for the system to be found in any many-body eigenstate with $m$ particles as a function of time. The empty markers are obtained by solving the quantum master equation whereas the solid lines are from the semiclassical description. We used $L=7$ and $g_{1}/g_{2}=\Gamma/g_{2}=0.1$.
The semiclassical state in Eq. (14) provides a complete description of the dynamics. For fermionic systems, the state satisfies Wick’s theorem [46], hence one can further rewrite it into a Gaussian form [47]:
$\hat{\rho}=\left(\det G\right)\exp\left[\sum_{i,j=0}^{L-1}K_{ij}(t)\hat{f}_{i}^{\dagger}\hat{f}_{j}\right],$ (16)
with the Gaussian kernel $K$ given by:
$K=\ln\left[(1-G)G^{-1}\right].$ (17)
[…]
$\chi_{k}(t)=\begin{cases}1-\exp(-\Gamma t/2)&{\rm for}\leavevmode\nobreak\ k<0,\\ 0&{\rm for}\leavevmode\nobreak\ k=0,\\ 1&{\rm for}\leavevmode\nobreak\ k>0.\end{cases}$ (18)
Finally, $G$ is related to $\chi$ by the same canonical transformation between $\hat{f}_{i}$ and $\hat{C}_{k}$, given in Eqs. (5) and (7). By using Eq. (18), it can be written as
$G_{ij}(t)=\sum_{k=1}^{(L-1)/2}\left[(1-e^{-\Gamma t/2})U_{-k,i}U_{-k,j}+U_{k,i}U_{k,j}\right],$ (19)
where $U_{\pm k,2i}=A_{ki}$ and $U_{\pm k,2i-1}=\pm B_{ki}$.
The Gaussian form (16) of the density matrix turns out to be extremely powerful when later calculating the entanglement content. We emphasize that for bosons the mixed state in Eq. (14) is not Gaussian anymore, despite having exactly the same form.
3.2 Entanglement measures
Quantifying quantum entanglement [48] is vital in quantum information theory, yet is highly non-trivial. Common entanglement measures include entanglement entropy [49], concurrence [50], and logarithmic negativity (or the closely related negativity) [51, 52]. However, entanglement entropy is applicable only to pure states, while concurrence is practical only for two-qubit systems, whether in pure or mixed states. For mixed states of more general systems, the logarithmic negativity (or negativity) has been widely used as one of very few practically computable entanglement measures. Nevertheless, as it requires finding the eigenvalues of the density matrix after partial transposition, its computational cost increases quickly with system size. To avoid complications arising from partial transposition, quantum mutual information is often used as well. Although mutual information is not an entanglement measure in the strict mathematical sense, it still captures many important aspects of quantum correlations, as well as classical correlations. Below, we will investigate the entanglement content of the mixed state (14) using the logarithmic negativity. However, as we illustrate in A, the universal features we find are observed in the mutual information as well.
The logarithmic negativity between two subsystems $A$ and $B$ is defined as [51]
$\mathcal{E}(A,B):=\log_{2}\left\|\hat{\rho}^{T_{A}}\right\|=\log_{2}\left\|\hat{\rho}^{T_{B}}\right\|,$ (20)
[…]
Recently, a new type of logarithmic negativity based on a partial time-reversal transformation has been suggested for fermionic systems [54, 55, 56]. Compared to the conventional (or bosonic) logarithmic negativity, this fermionic logarithmic negativity was shown to better capture the contributions to entanglement from Majorana fermions. However, the key advantage from our point of view lies in the fact that the partial time-reversal transformation of a fermionic Gaussian state remains Gaussian, allowing the fermionic logarithmic negativity to be calculated using a $2L\times 2L$ matrix of generalized Green’s functions [54]; that is, with a polynomial amount of resources and time. In our case, as the pairing correlations are zero, the fermionic logarithmic negativity is obtained directly from the $L\times L$ matrix of Green’s functions, see Eq. (19), and we could study half-SSH chains up to $L\sim 10^{3}$ without difficulties. For the bosonic case, the calculations are limited to relatively short chains.
3.3 Fermionic vs bosonic SSH models
Above, we have already made several remarks on technical differences between the bosonic and fermionic SSH models. Here, we summarize both fundamental and technical differences between them. It is remarkable that, despite these differences, the fermionic and bosonic models share the same universal dynamics of entanglement.
Both bosonic and fermionic systems in the half-SSH model exhibit a chiral symmetry, characterized by the existence of a unitary operator that anti-commutes with the Hamiltonian. This symmetry leads to a degenerate zero-energy subspace and a symmetric single-particle energy spectrum around zero energy. Additionally, the edge mode at zero energy reflects the non-trivial topological character of the band structure.
For large bosonic systems, the size of the Hilbert space poses significant computational challenges. Certain techniques, such as block diagonalization of $\hat{\rho}^{T_{A}}$, are employed to alleviate the computational burden. In contrast, fermionic systems can benefit from the fact that the relevant state is Gaussian, since the partial time-reversal transformation of Eq. (16) remains Gaussian.
In the next Section, we will present our results for both bosonic and fermionic systems in parallel, to highlight the common universal behavior despite the differences summarized above. We attribute the common features of entanglement dynamics to the chiral symmetry and topological properties, especially the zero-energy edge states.
4 Results
We now present our findings, which reveal intriguing dynamical phenomena and provide deeper insights into the relationship between dynamics and topological properties. These results contribute to a better understanding of how topological effects influence dynamical processes, and have significant implications for future developments of quantum technologies.
Figure 3: Universal entanglement revival, obtained by computing the logarithmic negativity ($\mathcal{E}(\{2\},\{0,1,3,\dots\})$ between the single-site subsystem $\{\hat{f}_{2}\}$ and the rest of the chain. Panels (a) and (b) consider the fermionic model at $g_{1}/g_{2}=0.1$ and $0.5$, respectively. Panels (c) and (d) are for the bosonic model, also using $g_{1}/g_{2}=0.1$ and 0.5. Other parameters are $g_{2}=1$ (arbitrary units) and $\Gamma/g_{2}=0.1$.
4.1 Universal entanglement revival
Figure 3 summarizes our main results, concerning the evolution of entanglement between the second even site and the rest of the system. The two left panels of Fig. 3 refer to the fermionic half-SSH chain, where we computed the fermionic logarithmic negativity for two values of $g_{1}$ deep in the topologically non-trivial phase ($g_{1}\ll g_{2}$). The two right panels of Fig. 3 show the conventional logarithmic negativity of the bosonic model, using the same set of parameters. In the fermionic case, we could calculate the fermionic logarithmic negativity for systems as large as $L\sim 10^{3}$, and three representative sizes are shown in panels (a) and (b) of Fig. 3. In the bosonic case, the calculations were limited to relatively small values of $L$, because of the computational difficulties discussed in Section 3, and panels (c) and (d) show the results for $L=3,5,\dots,11$. In both cases, two distinctive features are clearly seen: First, the entanglement revives after an initial decrease. Considering the very fragile nature of quantum entanglement in the presence of decoherence, an entanglement revival is unusual and has attracted much interest [16, 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 15, 14, 13]. Second, the temporal profile of the entanglement has a universal dependence, i.e., it is almost unaffected by the system size. Although in Fig. 3 we only show the entanglement between site $\hat{f}_{2}$ and the remaining sites, $\{\hat{f}_{0},\hat{f}_{1},\hat{f}_{3},\dots\hat{f}_{L}\}$, the presence of an universal revival behavior can be observed regardless of the specific choice of partitioning, as long as one subsystem is confined to the left edge. However, the visibility of the revival depends on the partitioning.
As demonstrated in a recent work [16], the entanglement revival behavior is due to a self-purification process directly related to the chiral symmetry and the associated zero-energy mode (not necessarily localized to edges): The system starts from a pure state with relatively large entanglement and, as it undergoes the decoherence process, decays to various intermediate states. Consequently, the purity of the state decreases significantly, which naturally leads to a considerable reduction in entanglement. After this intermediate stage, however, the system begins to recover its purity (and hence entanglement), because all particles have the fate of decaying out, except for the ones occupying the zero-energy edge mode. Eventually, the system stabilizes to the final state, i.e., the zero-energy mode, which is immune to the decay process. The entanglement in the steady state is always determined by this zero-energy mode.
While the above arguments explains the presence of the revival, its universal character is more intriguing, and is the main topic of the rest of the paper. In the following, we perform a detailed analysis of the size dependence, revealing that the universal behavior can be attributed to the topological feature of the system, especially the zero-energy edge mode.
4.2 Topological effects through the edge state
In Fig. 3, we have seen that the entanglement revival in the deep topological phase ($g_{1}/g_{2}\ll 1$) is universal, having a dynamical profile independent of system size. However, as the ratio $g_{1}/g_{2}$ increases to unity (i.e., the phase boundary of the topological phase is approached), the evolution of entanglement shows significant variations at small system sizes. When $g_{1}/g_{2}=0.9$, for example, the dynamical profile for $L=3$ is distinct from the profiles of larger systems ($L=19,1001$), as shown in the inset of Fig. 4. We find that the dynamical profiles remain universal as long as the system size is sufficiently large compared to a certain length scale, directly related to the localization length of the zero-energy edge mode. In the following, we examine this crossover scale in detail.
We define two length scales concerning the entanglement revival dynamics: The crossover length $L_{c}$ is defined as the minimum system size which, at a given ratio $g_{1}/g_{2}$, achieves a universal time dependence. In other words, the dynamical profiles of systems larger than $L_{c}$ are all universal (within a certain numerical tolerance $\epsilon$). More specifically, let $\mathcal{E}_{L}(t)$ be the logarithmic negativity for a system of length $L$. Then, choosing a sufficiently large $T$, at which the entanglement has reached its asymptotic value, we require that for all $L\geq L_{c}$
[…]
Numerically, we approximate the $L\to\infty$ limit by taking $L=1001$. We also choose $T=20$ and $\epsilon=0.05$. Figure 4 plots the crossover length $L_{c}$ (blue empty circles) as a function of $g_{1}/g_{2}$. It is important to note that $L_{c}$ closely resembles the localization length $\xi$ of the edge mode (black solid line). For instance, when $g_{1}/g_{2}=0.9$, the crossover length matches the localization length $L_{c}\approx\xi\approx 19$. This value is highlighted by a horizontal dashed line in the main panel of Fig. 4. In the inset, we can see that the detailed entanglement profile with $L=19$ follows quite faithfully the universal dynamics at $L=1001$.
Figure 4: Comparison of the crossover length $L_{c}$ (blue empty circles) and the effective length $L_{\mathrm{eff}}$ (red filled circles) with the localization length $\xi$ of the edge state (black solid line), as a function of the coupling ratio $g_{1}/g_{2}$. The red dashed line marks the values $L_{c}\approx 19$, computed at $g_{1}/g_{2}=0.9$. Inset: time dependence of the logarithmic negativity between the single-site subsystem $\{\hat{f}_{2}\}$ and the rest of the chain. Here we take $g_{1}/g_{2}=0.9$ and $L=3$, $19$, $1001$ (from top to bottom).
[…]
The explicit dependence of $L_{\rm eff}$ on $g_{1}/g_{2}$ is shown in Fig. 4. As seen, also the effective length $L_{\mathrm{eff}}$ (red filled circles) agrees well with the localization length $\xi$ of the edge mode. That both $L_{c}$ and $L_{\rm eff}$ follow closely the localization length $\xi$ strongly suggest that the spatial range of the edge mode is a critical factor in determining the onset of the universal entanglement revival behavior.
Figure 5: The visibility of entanglement revival as a function of the coupling ratio $g_{1}/g_{2}$, for $L=1001$ and $\Gamma/g_{2}=0.1$.
Another feature that supports the topological origin of the entanglement revival is the visibility, defined as the difference between the asymptotic value ($t\to\infty$) and the minimum of the logarithmic negativity. The details of the visibility differ between the fermionic and bosonic cases, and they are also affected by system parameters. However, as illustrated in Fig. 5 for a long fermionic chain, we find the following generic qualitative features: The visibility is zero at $g_{1}/g_{2}=0$ and first increases with $g_{1}/g_{2}$, reaching a maximum around $g_{1}/g_{2}\approx 0.5$. Then, it sharply drops to zero (i.e., the entanglement revival disappears) as $g_{1}/g_{2}$ approaches 1. The visibility vanishing at $g_{1}/g_{2}=0$ is trivial, since the unit cells are decoupled. However, the disappearance of the entanglement revival when $g_{1}/g_{2}\to 1$, which is the boundary of the topological phase, strongly suggests that the entanglement revival is closely connected with the edge mode.
Finally, while in this section we have exclusively discussed the logarithmic negativity, the universal revival can also be found in the mutual information (see A). The von Neumann and Rényi entropy densities exhibit a size-independent behavior as well (see B), suggesting that the topological properties of the system influence more generally the dynamical evolution and robustness of local correlations, both quantum and classical.
5 Conclusion
In this study of the fermionic and bosonic SSH models, characterized by the presence of a chiral symmetry, we explored the dynamics of entanglement of a single site with the rest of the chain, in the presence of decoherence which preserves the chiral symmetry. Our results reveal a remarkable entanglement revival which, for a fixed coupling ratio, acquires a universal profile when the system length exceeds the localization length of the edge states. Therefore, the effective length scale for the onset of universality can be identified with the localization length. These findings highlight a direct connection between the universal entanglement revival and the topological phase.
Acknowledgments
C.D. and M.-S.C. have been supported by the National Research Function (NRF) of Korea (Grant Nos. 2022M3H3A106307411 and 2023R1A2C1005588) and by the Ministry of Education through the BK21 Four program. S.C. acknowledges support from the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0301602) and the National Science Association Funds (Grant No. U2230402).
Appendix A Mutual information
The mutual information is a measure of total correlations, both classical and quantum, between two subsystems. It provides a different perspective on the state of the system than entanglement measures, like the logarithmic negativity. Furthermore, the mutual information is usually easier to compute than most entanglement measures, making it a practical alternative for larger systems. Here, we show that the mutual information, like the logarithmic negativity, exhibits a universal entanglement revival, which indicates that universality is not limited to a specific entanglement measure.
The mutual information between subsystems $A$ and $B$ is defined as:
$\mathrm{I}_{\gamma}(A,B):=S_{\gamma}(\hat{\rho}_{A})+S_{\gamma}(\hat{\rho}_{B})-S_{\gamma}(\hat{\rho}),$ (22)
where $\hat{\rho}_{A/B}$ is the reduced density matrix of subsystem $A/B$ and $\hat{\rho}$ the full density matrix, while the Rényi entropy of order $\gamma$ is defined as:
$S_{\gamma}(\hat{\rho}):=\frac{1}{1-\gamma}\log_{2}\mathrm{Tr}(\hat{\rho}^{\gamma}).$ (23)
In the limit $\gamma\rightarrow 1$, the Rényi entropy becomes the von Neumann entropy, $S_{\mathrm{vN}}(\hat{\rho})=-\mathrm{Tr}(\hat{\rho}\log_{2}\hat{\rho})$. However, computing the von Neumann entropy can still be computationally intensive for large systems. In the case with $\gamma=2$, the Rényi entropy simplifies to the second-order Rényi entropy, $S_{2}(\hat{\rho})=-\log_{2}\mathrm{Tr}(\hat{\rho}^{2})$, which is easier to compute and advantageous for large or complex systems.
Figure A1: Time dependence of the von Neumann mutual information in the fermionic model. $\mathrm{I}_{\mathrm{vN}}(\{2\},\{0,1,3,...\})$ indicates the von Neumann mutual information between the single-site subsystem $\{\hat{f}_{2}\}$ and the rest of the chain. In panels (a) and (b) we use $g_{1}/g_{2}=0.1$ and $0.5$, respectively. In both panels, $\Gamma/g_{2}=0.1$.
For fermionic Gaussian states, the von Neumann entropy can be calculated efficiently from the correlation function [57]. In Fig. A1, we can observe that the entanglement revival might disappear in certain limits, such as $g_{1}\ll g_{2}$, because the mutual information is not a proper measure of quantum entanglement. However, the universal behavior always persists, which is in agreement with the results for the logarithmic negativity presented in the main text. In addition, both measures suggest that, for a fixed coupling ratio $g_{1}/g_{2}$, the universal behavior can be observed when the size of the half-SSH model exceeds a certain crossover length. The consistency in behavior between the two measures underscores the robustness of the universal entanglement revival phenomenon, confirming once more that the universal behavior is determined by the inherent properties of the system.
Figure A2: The revival behavior of von Neumann mutual information in the bosonic model for coupling ratios (a) $g_{1}/g_{2}=0.1$ and (b) $g_{1}/g_{2}=0.5$. Here, we set $\Gamma/g_{2}=0.1$.
For the bosonic case, computing the von Neumann mutual information is more challenging than fermions, since the Hilbert space still grows exponentially with the chain length, increasing the difficulty in computing eigenvalues. Currently, we are able to calculate the von Neumann mutual information up to $L=13$. Figure A2 illustrates the entanglement dynamics for bosonic systems with different coupling ratios and chain lengths. As seen, despite the maximum length being much shorter than for fermions, the universal behavior is still present. Thus, the conclusions for bosons are very similar to those for fermions in Fig. A1, and to the behavior of the logarithmic negativity.
Due to the limitations in calculating the von Neumann mutual information for bosonic systems, we extended our analysis to the the Rényi mutual information, which can be obtained for larger system sizes (specifically, we consider $L=11,15,19$, and $23$). As shown in Fig. A3, also the Rényi mutual information displays a universal revival. The larger accessible system sizes allow us to obtain good convergence also in Fig. A3(b), where a larger coupling ratio $g_{1}/g_{2}=0.5$ is assumed. This extension further confirms that the universal behavior is not restricted to the von Neumann mutual information and logarithmic negativity, offering new evidence of its robustness in larger bosonic systems.
Figure A3: Revival behavior of the Rényi mutual information in the bosonic model, for coupling ratios (a) $g_{1}/g_{2}=0.1$ and (b) $g_{1}/g_{2}=0.5$. We denote the Rényi mutual information between the single-site subsystem $\{\hat{f}_{2}\}$ and the rest of the chain by $\mathrm{I}_{2}(\{2\},\{0,1,3,\cdots\})$. The inset of panel (b) is a zoom-in of the main plot. We have used $\Gamma/g_{2}=0.1$.
Appendix B Entropy density
We also consider the average entropy $s_{\gamma}(\hat{\rho})=S_{\gamma}(\hat{\rho})/n$, where $S_{\gamma}(\hat{\rho})$ is defined in Eq. (23) and $n=(L-1)/2$ is the number of two-site unit cells. These quantities can be easily evaluated within the semiclassical approximation of $\hat{\rho}$, given by Eqs. (14) and (15). In particular, the von Neumann entropy density $s_{\text{vN}}(\rho)$ is obtained as follows:
$s_{\mathrm{vN}}=-\log_{2}(1-e^{-\Gamma t/2})+e^{-\Gamma t/2}\log_{2}(-1+e^{\Gamma t/2}),$ (24)
and the Rényi entropy density of order 2 reads:
$s_{2}=-\log_{2}(1+2e^{-\Gamma t}-2e^{-\Gamma t/2}).$ (25)
The qualitative time dependence of these quantities differs from the logarithmic negativity, since the system is in a pure state at $t=0$ and $t\to\infty$ (when the entropy is zero) and there is no revival behavior. Instead, the entropy density attains its maximum value $s_{\mathrm{vN}}=s_{2}=1$ at $\Gamma t=2\ln 2$. The increase of the entropy at intermediate times in in qualitative agreement with the reduction of purity and the occurrence of a minimum in the logarithmic negativity. However, in contrast to the entanglement, the maximum entropy always occurs at $\Gamma t=2\ln 2$, independent of the coupling strengths $g_{1,2}$. This difference is due to the fact that the probabilities of the eigenstates, given in Eq. (15), do not depend on $g_{1,2}$. Instead, the entanglement is obviously affected by the specific eigenstates which diagonalize $\hat{\rho}$, thus the position of the minimum depends the coupling ratio $g_{1}/g_{2}$. Despite these remarks, we note that both Eqs. (24) and (25) are independent on $L$, i.e., they display a universal behavior, demonstrating that local quantum correlations are robust with respect to the overall system size. These results suggest that the robustness of local quantum correlations is likely due to the topological properties of the SSH model, contributing to the stability of the entanglement revival under different conditions.